OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system −3 3

OpenStudy (zzr0ck3r):

multiply the first one by -4/3 and we get -16/3x+4y = -4 add that to 5x-4y = 3 (-16/3) x + 5x = -1 x(-16/3+5) = 1 x(-1/3) = 1 x = -3

OpenStudy (anonymous):

thank you but what did the / mean so i can write it in my notes

OpenStudy (zzr0ck3r):

a/b = \[\frac{ a }{ b }\]